# AB-UPT: Scaling Neural CFD Surrogates for High-Fidelity Automotive Aerodynamics Simulations via Anchored-Branched Universal Physics Transformers

### v2

@@ A.3 Analysis: Anchor Attention
Table 9: Analysis of the effectiveness of anchor attention vs a simple $k$-NN interpolation. Anchor attention can learn more complex interactions than a simple interpolation.
@@
$\mathbf{p}_{s}$ $\mathbf{u}$
k-NN interpolation 0.00685 0.06643
Anchor-attention 0.00260 0.00684
@@ B.1 Datasets
ShapeNet Car
The ShapeNet Car dataset, a subset of the ShapeNet dataset (Chang et al., 2015) as introduced by Umetani & Bickel (2018), specifically focuses on the car-labeled data points. Each simulation in this dataset includes a surface mesh with 3,682 points and a corresponding volume mesh with 28,504 points. Following (Alkin et al., 2024a), we remove outlier points that do not belong to the surface mesh (which results in a total of 3,586 points per surface mesh). As model input, we only consider the points on the surface mesh (i.e., without using additional signed distance function values or surface normals). Our models are then trained to predict pressure values on the surface mesh and velocity values within the volume. We adopt the same training and testing split as in Wu et al. (2024), which reserves param0 of the dataset for testing. This results in 789 samples for training and 100 samples for testing.
@@ B.2 Evaluation metrics
Our dataset consists of a set of $M$ target vectors in point clouds $\mathcal{Y}=\{\mathbf{Y}^{(1)},\mathbf{Y}^{(2)},\dots\mathbf{Y}^{(M)}\}$ and corresponding predictions $\hat{\mathcal{Y}}=\{\hat{\mathbf{Y}^{(1)}},\hat{\mathbf{Y}^{(2)}},\dots\hat{\mathbf{Y}^{(M)}}\}$ where the average realtive L2 error is defined as:
@@ B.3 AB-UPT
In the following subsection, we use the a/b/c notation to indicate hyperparameters for ShapeNet-Car/AhmedML/DrivAerML, respectively. We use a radius of $9/0.25/0.25$ for the supernode pooling in the geometry branch. The geometry branch consists of a supernode pooling layer, followed by a single transformer block with self-attention. The number of supernodes in the geometry branch is $3586/16384/16384$, which we also use as the number of surface anchor points in the surface branch. The number of volume anchor points is $4096/16384/16384$. When comparing against baselines, we train without any query tokens, i.e., we create predictions for each anchor token and calculate the loss those. In other settings (e.g., when training with the CAD geometry), we also use query tokens during training, where we set the number of query tokens to be equal to the number of anchor tokens.
We choose the architecture such that surface/volume tokens traverse $12$ transformer blocks. The first block in each branch is a cross-attention block to the output of the geometry branch. The next $8/8/4$ blocks consist of interleaved cross-attention and self-attention blocks, where cross-attention blocks exchange information between branches by using tokens of the other branch as keys and values. The self-attention blocks operate only within their branch. The last $3/3/7$ blocks are also self-attention blocks. Afterwards, a linear layer decodes the token representation into 4 surface variables (1 pressure dimension, $3$ wallshearstress dimensions) and $7$ volume variables ($1$ pressure dimension, 3 velocity dimension, $3$ wallshearstress dimensions).
We use a standard ViT (Dosovitskiy et al., 2021)-style pre-norm transformer blocks where we additionally add Rotary Position Embeddings (RoPE(Su et al., 2024) to all blocks.
We train models for 500 epochs using batch size 1, LION optimizer (Chen et al., 2023) with peak learning rate 5e-5, weight decay 5e-2, a linear warmup for 5% of the training duration, followed by cosine decay with end learning rate of 1e-6. We train in float16 mixed precision and employ a mean squared error. Training takes roughly $3/7/7$ hours on a single NVIDIA H100 GPU and occupies 4GB of GPU memory. Similar to (Wu et al., 2024), we choose batch size 1 because it obtains the best accuracy.
@@ B.4 Baseline models
Transolver (Wu et al., 2024) is a transformer-based baseline model and, at the time of writing, the state-of-the-art on ShapeNet Car. It introduces the Physics-Attention mechanism, where each layer in the Transolver model takes a finite discrete point cloud representation of an input geometry as input, and maps each point to a learnable slice (also referred to as physics tokens). Points with similar physical properties are mapped to the same slice. First, each surface mesh point is mapped to slice weights, which indicate the degree to which each point belongs to a slice. Next, the slice weights are used to aggregate point features into physics-aware tokens. Multi-head self-attention is applied to these physics-aware tokens, rather than directly to the input points, which reduces the computational cost of the self-attention layer. Finally, after the self-attention layer, the physics-aware tokens are transformed back to mesh input points by deslicing. Afterward, a feed-forward layer is applied to the individual input point representations.
@@
As Transolver++ (Luo et al., 2025) has no publicly available implementation, we attempted to reimplement it to the best of our abilities and tried to infer implementation details that are not outlined in the paper. However, contrary to the results of the original paper, our reimplementation performed consistently worse than the vanilla Transolver (Wu et al., 2024). Therefore, we do not consider it in the experiments section.
@@ B.4.1 Data pre-processing
. For Shapenet, we train our models with 3586 points from the surface mesh and 4096 points in the volume. For both AhmedML and DrivAerML, we train our models with 16384 surface and volume points. To make the comparison with baseline models possible in Section 4.4, we chunk the evaluation meshes into chunks of 16384 for both the surface and volume.
@@ B.5.1 General configurations
For each baseline model, we used the AdamW (Loshchilov & Hutter, 2019) optimizer, with the exception of AB-UPT/UPT/Transformer, where we use LION (Chen et al., 2023). We tuned the learning rate of models trained with AdamW by sweeping over $\{1e-3,5e-3,2e-3,5e-3,1e-4,5e-5.2e-5\}$ to achieve optimal performance per baseline. Whenever LION is used, we use 5e-5 as this value has performed consistently well. Training was conducted using either float16 or bfloat16 precision, depending on which yielded the best results. An exception was made for GINO, as it exhibited unstable behavior with float16/bfloat16 mixed precision. All models were trained for $500$ epochs with a batch size of $1$. We applied a weight decay of $0.05$ and a gradient clipping value of $1$. A cosine learning rate schedule was implemented, including a $5$% warm-up phase and a final learning rate of $61.0e-$ 6
@@ B.5.2 Model specific hyper-parameters
Table 10: Training hyper-parameters and the number of trainable parameters for each model. Where a parameter varies across datasets, it’s presented in the format ShapeNetCar, AhmedML, DrivAerML; otherwise, a single value applies to all.
@@
GRAPH U-NET $1e^{-3}$float16, bfloat16, float16} - 14.11M
@@
Transolver $1e^{-3}$ bfloat16 - 5.49M
Transformer $5e^{-5}$ float16 - 5.49M
AB-UPT $5e^{-5}$ float16 - {6.71M, 8.75M, 8.75M}
@@ B.5.3 Evaluation variance reduction
    each test sample is divided into chunks to match the number of inputs used during training, where the number of volume cells is first subsampled to match the number of surface cells.
@@ Appendix C Baseline fairness and reproducibility
Table 11: Comparison of our baseline model evaluated on ShapeNetCar, run in the same experimental set-up as in the Transolver paper. Evaluating median and best performance over 5 training runs with different seeds for $\mathbf{p}_{s}$ and $\mathbf{u}$, alongside the results reported in the Transolver paper. The results for AB-UPT are from Table 4 (*).
@@
Model $\mathbf{p}_{s}$ $\mathbf{u}$ $\mathbf{p}_{s}$ $\mathbf{u}$ $\mathbf{p}_{s}$ $\mathbf{u}$
@@
In this work, we train and evaluate AB-UPT and all the baseline models in our unified experimental framework. Some of our baselines are taken from the Neural-Solver-Library (Wu et al., 2024). To validate the correctness of our experimental framework and baseline implementations, we rerun all the ShapeNetCar experiments with the experimental setup as reported in Wu et al. (2024) (including the same input features: SDF and normal vectors in our own framework. We report the median/best pressure ($\mathbf{p}_{s}$) and velocity ($\mathbf{u}$) values over five training runs with different seeds, and where possible, we compare the values with the evaluation metrics reported in Wu et al. (2024). For each Transformer-based model, we use $8$ Transformer blocks, with $8$ attention heads, and a hidden dimensionality of $256$, similar to Transolver.

### v3

@@ A.6 Anchor attention versus k-NN interpolation
Table 13: Analysis of the effectiveness of anchor attention vs a simple k-NN interpolation. Anchor attention can learn more complex interactions than a simple interpolation. Values denote test MSE losses on AhmedML.
@@
$\bm{p}_{s}$ $\bm{u}$ $\bm{\omega}$
k-NN interpolation 0.00703 0.06641 0.5748
Anchor attention 0.00276 0.00657 0.0397
A.7 Extended comparison against Transolver
In Transolver (Wu et al., 2024), the quadratic attention is limited to a subset of tokens by slicing the input domain into a reduced domain (i.e., physics tokens). Throughout the experiments in the main paper, we use 512 slices. In this section, we investigate the impact of this hyperparameter in terms of performance (test loss), memory consumption, and runtime. Figure 11 shows that increasing the number of slices slightly improves test loss at the cost of additional memory and runtime requirements. As the slicing mechanism is implemented via slow and memory-heavy torch.einsum operations (following the original implementation [^4] https://github.com/thuml/Transolver/blob/main/Car-Design-ShapeNetCar/models/Transolver.py), it requires much more memory than AB-UPT, which leverages the constant memory complexity of FlashAttention (Dao et al., 2022) (which results in constant memory complexity for the attention layer, but the overall memory complexity is still linear). For larger problem sizes (e.g., DrivAerML), the memory bottleneck becomes significant and requires memory-saving techniques such as activation checkpointing to fit onto a single GPU, which in turn increases runtime. Note that we improved speed and memory consumption of the original Transolver implementation by using FlashAttention for the self-attention between slice tokens and employing mixed-precision training. Further optimizations (e.g., specialized hardware implementations of the transolver attention) are certainly possible but are beyond the scope of this work. Note that we also do not optimize AB-UPT beyond the use of FlashAttention555torch.nn.functional.scaled_dot_product_attention using PyTorch 2.4.1 with CUDA 12.4 with mixed-precision training. Also AB-UPT could be optimized further, e.g., by employing FlashAttention-3 (Shah et al., 2024).
Figure 11: Performance (left), memory consumption (center) and runtime (right) on ShapeNet-Car (top) and DrivAerML (bottom) of transolver with variable number of slices (i.e., latent tokens). Increasing the number of slices slightly improves performance, but does not come close to AB-UPT. Additionally, increasing the number of slices greatly increases memory and compute demands. The number of surface/volume points for this study is 4K/4K for ShapeNet-Car and 65K/65K for DrivAerML.
A.8 Geometry branch design
We evaluate various design choices of the geometry branch in Table 11(f).
  1. (a)
    Subtracting the position of the supernode before encoding the positions of the incoming nodes allows the model to develop local filters instead of learning the same filter for every location in the domain. Although performance is relatively similar, we choose relative positions to enforce learning local patterns in the message passing of the supernode pooling layer. Note that we concatenate the absolute position embedding after supernode pooling. This is conceptually similar to Vision Transformers (Dosovitskiy et al., 2021) where pixels are encoded into patches using local operations without absolute positions embedding. Afterwards, absolute positions are integrated into patch tokens.
  2. (b)
    Increasing the number of geometry points, i.e., the resolution of the geometry input, does not significantly improve performance while increasing runtime.
  3. (c)
    Connecting points to supernodes via a k-NN graph can obtain similar results to a radius graph. However, using a k-NN graph makes the model dependent on the number of geometry inputs as more points would reduce the distance to the furthest neighbor. The radius graph does not have this disadvantage as it is similar to Graph Neural Operators (Li et al., 2020).
  4. (d)
    Even high-resolution geometries like the DrivAerML shapes can be captured with a relatively low number of geometry inputs and geometry tokens.
  5. (e)
    Integrating the geometry information directly at the start of the model shows better results than integrating it later on or attending multiple times to the geometry tokens via multiple cross-attention blocks in the surface/volume branches.
  6. (f)
    The geometry branch benefits from local aggregation, i.e., supernode pooling, followed by a global information exchange, i.e., a Transformer block. Adding more Transformer blocks is not necessary. Note that using neither supernode pooling nor Transformer blocks corresponds to not using the geometry branch at all where the cross-attention block is replaced by a self-attention block to keep the number of parameters of the surface/volume branch constant. Therefore, showing the benefits of using a geometry branch.
We note that the slight performance differences of (a-d) are negligible, and we choose designs/hyperparameters based on conceptual motivations (a, c) or lower runtimes (b, c).
Table 14: Geometry branch design study on DrivAerML. Test losses are multiplied by 10. Default settings
Test loss
Absolute in domain 1.360
Relative to supernode 1.357
#Points Test loss
64K 1.357
128K 1.357
Test loss
Radius graph (r=0.25) 1.357
k-NN graph (k=16) 1.361
k-NN graph (k=32) 1.362
#Points #Tokens Test loss
64K 16K 1.357
64K 32K 1.355
128K 32K 1.360
Location Test loss
Start 1.357
Mid 1.425
End 1.395
Start + mid 1.390
Start + end 1.360
Supernode pooling #Blocks Test loss
0 1.405
0 1.370
1 1.357
2 1.366
4 1.364
@@ B.1 Datasets
ShapeNet-Car
The ShapeNet-Car dataset, a subset of the ShapeNet dataset (Chang et al., 2015) as introduced by Umetani & Bickel (2018), specifically focuses on the car-labeled data points. Each simulation in this dataset includes a surface mesh with 3,682 points and a corresponding volume mesh with 28,504 points. Following (Alkin et al., 2024a), we remove outlier points that do not belong to the surface mesh (which results in a total of 3,586 points per surface mesh). As model input, we only consider the points on the surface mesh (i.e., without using additional signed distance function values or surface normals). Our models are then trained to predict pressure values on the surface mesh and velocity values within the volume. We adopt the same training and testing split as in Wu et al. (2024), which reserves param0 of the dataset for testing. This results in 789 samples for training and 100 samples for testing.
@@ B.2 Evaluation metrics
Our dataset consists of a set of $M$ target vectors in point clouds $\mathcal{Y}=\{\mathbf{Y}^{(1)},\mathbf{Y}^{(2)},\dots\mathbf{Y}^{(M)}\}$ and corresponding predictions $\hat{\mathcal{Y}}=\{\hat{\mathbf{Y}}^{(1)},\hat{\mathbf{Y}}^{(2)},\dots\hat{\mathbf{Y}}^{(M)}\}$ where the average relative L2 error is defined as:
@@ B.3 AB-UPT
In the following subsection, we use the a/b/c notation to indicate hyperparameters for ShapeNet-Car/AhmedML/DrivAerML, respectively. We use a radius of $9/0.25/0.25$ for the supernode pooling in the geometry branch. The geometry branch consists of a supernode pooling layer, followed by a single Transformer block with self-attention. The number of supernodes in the geometry branch is $3586/16384/16384$, which we also use as the number of surface anchor points in the surface branch. The number of volume anchor points is $4096/16384/16384$. When comparing against baselines, we train without any query tokens, i.e., we create predictions for each anchor token and calculate the loss for those. In other settings (e.g., when training with the CAD geometry), we also use query tokens during training, where we set the number of query tokens to be equal to the number of anchor tokens.
We choose the architecture such that surface/volume tokens traverse $12$ Transformer blocks. The first block in each branch is a cross-attention block to the output of the geometry branch. The next $8/8/4$ blocks consist of interleaved cross-attention and self-attention blocks, where cross-attention blocks exchange information between branches by using tokens of the other branch as keys and values. The self-attention blocks operate only within their branch. The last $3/3/7$ blocks are also self-attention blocks. Afterwards, a linear layer decodes the token representation into 4 surface variables (1 pressure dimension, $3$ wall shear stress dimensions) and $7$ volume variables ($1$ pressure dimension, 3 velocity dimension, $3$ vorticity dimensions).
We use a standard ViT (Dosovitskiy et al., 2021)-style pre-norm Transformer blocks where we additionally add Rotary Position Embeddings (RoPE(Su et al., 2024) to all blocks. For the message passing of the supernode pooling layer, we the positions relative to the supernode and the magnitude of the distance as input to the message passing. This is then followed by embedding the position of the supernode, concatenating it with the output of the message passing and down-projecting it to the original dimension in order to integrate absolute positions after the message passing. This is in contrast to Alkin et al. (2024a), where the concatenation of the positions of the two connected nodes are used directly for message passing.
We train models for 500 epochs using batch size 1, LION optimizer (Chen et al., 2023) with peak learning rate 5e-5, weight decay 5e-2, a linear warmup for 5% of the training duration, followed by cosine decay with end learning rate of 1e-6. We train in float16 mixed precision and employ a mean squared error. Training takes roughly $3/7/7$ hours on a single NVIDIA H100 GPU and occupies 4GB of GPU memory. Similar to Wu et al. (2024), we choose batch size 1 because it obtains the best accuracy.
@@ B.4 Baseline models
Transolver (Wu et al., 2024) is a transformer-based baseline model and, at the time of writing, the state-of-the-art on ShapeNet-Car. It introduces the Physics-Attention mechanism, where each layer in the Transolver model takes a finite discrete point cloud representation of an input geometry as input, and maps each point to a learnable slice (also referred to as physics tokens). Points with similar physical properties are mapped to the same slice. First, each surface mesh point is mapped to slice weights, which indicate the degree to which each point belongs to a slice. Next, the slice weights are used to aggregate point features into physics-aware tokens. Multi-head self-attention is applied to these physics-aware tokens, rather than directly to the input points, which reduces the computational cost of the self-attention layer. Finally, after the self-attention layer, the physics-aware tokens are transformed back to mesh input points by deslicing. Afterward, a feed-forward layer is applied to the individual input point representations.
@@
As Transolver++ (Luo et al., 2025) has no publicly available implementation, we attempted to reimplement it to the best of our abilities and tried to infer implementation details that are not outlined in the paper. However, contrary to the results of the original paper, our reimplementation performed consistently worse than the vanilla Transolver (Wu et al., 2024). Therefore, we do not consider it in the experiments section. We refer to Appendix C.2 for the full reproducibility ablation.
@@ B.4.1 Data pre-processing
For Shapenet-Car, we train our models with 3586 points from the surface mesh and 4096 points in the volume. For both AhmedML and DrivAerML, we train our models with 16384 surface and volume points. To make the comparison with baseline models possible in Section 4.4, we chunk the evaluation meshes into chunks of 16384 for both the surface and volume.
@@ B.5.1 General configurations
For each baseline model, we used the AdamW (Loshchilov & Hutter, 2019) optimizer, with the exception of AB-UPT/UPT/Transformer, where we use LION (Chen et al., 2023). We tuned the learning rate of models trained with AdamW by sweeping over $\{1e^{-3},5e^{-3},2e^{-3},5e^{-3},1e^{-4},5e^{-5},2e^{-5}\}$ to achieve optimal performance per baseline. Whenever LION is used, we use $5e^{-5}$ as this value has performed consistently well. Training was conducted using either float16 or bfloat16 precision, depending on which yielded the best results. An exception was made for GINO, as it exhibited unstable behavior with float16/bfloat16 mixed precision. All models were trained for $500$ epochs with a batch size of $1$. We applied a weight decay of $0.05$ and a gradient clipping value of $1$. A cosine learning rate schedule was implemented, including a $5$% warm-up phase and a final learning rate of $1e^{-6}$
@@ B.5.2 Model specific hyper-parameters
Table 15: Training hyper-parameters and the number of trainable parameters for each model. Where a parameter varies across datasets, it’s presented in the format ShapeNet-Car, AhmedML, DrivAerML; otherwise, a single value applies to all.
@@
GRAPH U-NET $1e^{-3}$float16, bfloat16, float16} - 14.1M
@@
Transolver $1e^{-3}$ bfloat16 - 5.5M
Transformer $5e^{-5}$ float16 - 5.5M
AB-UPT $5e^{-5}$ float16 - {6.7M, 6.7M, 8.8M}
@@ B.5.3 Evaluation variance reduction
    Each test sample is divided into chunks to match the number of inputs used during training, where the number of volume cells is first subsampled to match the number of surface cells. All chunks of a sample are concatenated before calculating metrics.
@@ Appendix C Baseline fairness and reproducibility
C.1 Transolver and baseline reproducibility
Table 16: Comparison of our baseline model evaluated on ShapeNet-Car, run in the same experimental set-up as in the Transolver paper. Evaluating median and best performance over 5 training runs with different seeds for $\bm{p}_{s}$ and $\bm{u}$, alongside the results reported in the Transolver paper. The results for AB-UPT are from Table 4 (*).
@@ C.1 Transolver and baseline reproducibility
Model $\bm{p}_{s}$ $\bm{u}$ $\bm{p}_{s}$ $\bm{u}$ $\bm{p}_{s}$ $\bm{u}$
@@
In this work, we train and evaluate AB-UPT and all the baseline models in our unified experimental framework. Some of our baselines are taken from the Neural-Solver-Library (Wu et al., 2024). To validate the correctness of our experimental framework and baseline implementations, we rerun all the ShapeNet-Car experiments with the experimental setup (i.e., data loading, training hyper-parameters, model parameters) as reported in Wu et al. (2024) (including the same input features: SDF, input coordinates, and normal vectors) in our own framework. We report the median/best pressure ($\bm{p}_{s}$) and velocity ($\bm{u}$) values over five training runs with different seeds, and where possible, we compare the values with the evaluation metrics reported in Wu et al. (2024). For each Transformer-based model, we use $8$ Transformer blocks, with $8$ attention heads, and a hidden dimensionality of $256$, similar to Transolver.
@@
C.2 Reproducibility of Transolver++
Table 17: Attempt to implemented Transolver++ on ShapeNet-Car. We started with the Transolver model, and step-by-step tuned the implementation into Transolver++ by: (i) adding the network reparameterization of the slicing network, (ii) using the Gumbel-Softmax instead of the softmax function, and (iii) adding the adaptive temperature parameter. We report evaluation metrics (surface pressure $\bm{p}_{s}$ and volume velocity $\bm{u}$) for a single training run.
L2 error ($\downarrow$)
$\bm{p}_{s}$ $\bm{u}$
Transolver 6.48 1.57
+ reparameterization 6.72 1.63
+ Gumbel-Softmax 7.28 1.63
Transolver++ + adaptive temperature 7.57 1.76
Transolver++ (Luo et al., 2025) is the parallelizable, more efficient, and accurate successor to Transolver (Wu et al., 2024). Since the code for Transolver++ is not publicly available at the time of writing, we attempted to reimplement the model ourselves. Our goal was to include it in our baseline benchmark experiments, as shown in Table 4. Turing Transolver into Transolver++ involves three steps (excluding the parallelization)
  1. (i)
    Removing the over-parameterization of the slicing (+ reparameterization);
  2. (ii)
    Use a Gumbel-Softmax instead of a softmax (+ Gumbel-Softmax);
  3. (iii)
    Use adaptive temperature parameters (+ adaptive temperature).
In Table 17 we report the L2 error for the surface pressure and volume velocity for Transolver and the results after applying each individual step sequentially. Unfortunately, all the additions that define Transolver++ consistently led to worse evaluation performance. Therefore, we decided to exclude Transolver++ as a benchmark model in this work.